OpenStudy (anonymous):

A multiple choice test has 10 questions. Each question has four answer choices. a. What is the probability a student randomly guesses the answers and gets exactly six questions correct? b.Is getting exactly 10 questions correct the same probability as getting exactly zero correct? c. Describe the steps needed to calculate the probability of getting at least six questions correct if the student randomly guesses.

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

OK... have you gone into binomeal expansion?

OpenStudy (anonymous):

Yeah a little.

OpenStudy (e.mccormick):

Guess what this will need...

OpenStudy (anonymous):

binomial expansion...?

OpenStudy (e.mccormick):

Yah...

OpenStudy (e.mccormick):

So if your sucess ia 1/4 and the falure is 3/4

OpenStudy (anonymous):

(p+q)?

OpenStudy (e.mccormick):

Yes, that is what you could set those two. Values like that, then find where your specific case falls, and expand that.

OpenStudy (anonymous):

is it ^8?

OpenStudy (e.mccormick):

^8? For what?

OpenStudy (anonymous):

(p+q)^8

OpenStudy (e.mccormick):

There are 10 questions, \((p+q)^{10}\)

OpenStudy (anonymous):

Oh duh. So can I just use Pascal's triangle to expand it?

OpenStudy (e.mccormick):

Yes, if you have the triangle that deep, or the theorem for binomial expansion.

OpenStudy (anonymous):

whats the theorem?

OpenStudy (e.mccormick):

\[(p+q)^n=\sum_{k=0}^n\left( \begin{matrix}n\\k\end{matrix}\right)p^{n-k} q^k\]

OpenStudy (e.mccormick):

Where \[\left( \begin{matrix}n\\k\end{matrix}\right)=\frac{n!}{(n-k)!k!}\]

OpenStudy (anonymous):

o.O I am so confused.

OpenStudy (e.mccormick):

For low values of n, Pascal's Triangle is better. But if say n=200 and you want the 123rd term, the formula is better.

OpenStudy (anonymous):

okay so it would be (p+q)^10 on the left but what is k in the formula?

OpenStudy (e.mccormick):

It is a start point in the sigma sum.

OpenStudy (e.mccormick):

If you have not seen sigma, \(\Sigma\) notation, stick with Pascal.

OpenStudy (anonymous):

I have seen it. would k=1?

OpenStudy (e.mccormick):

It starts at 0 and goes to n.

OpenStudy (e.mccormick):

Know how in any line of pascals, the first term is \(x^n\) and the second is \(x^{n-1}y\)? That is because k starts at 0. \(x^n=x^ny^0\) and \(x^{n-1}y=x^{n-1}y^1\)?

OpenStudy (anonymous):

so far I have \[{p+q}^2 =\sum_{k=0}^{10}\] right?

OpenStudy (e.mccormick):

\[(p+q)^{10} =\sum_{k=0}^{10}\]

OpenStudy (anonymous):

right. so it should be \[(p+q)^{10}=\sum_{k=0}^{10} (\frac{ 10! }{ (10)!0! }) p ^{10} q ^{0}\]

OpenStudy (e.mccormick):

well, uhhh... that would be... sort of all expanded for the 10th spot. Let me back it up a step.

OpenStudy (e.mccormick):

\[(p+q)^10=\sum_{k=0}^10\left(\frac{n!}{(n-k)!k!}\right)p^{n-k} q^k\]

OpenStudy (e.mccormick):

Oops! (p+q)^{10}=\sum_{k=0}^{10}\left(\frac{n!}{(n-k)!k!}\right)p^{n-k} q^k

OpenStudy (e.mccormick):

Aaargh! Forgot my braces. LOL \[(p+q)^{10}=\sum_{k=0}^{10}\left(\frac{n!}{(n-k)!k!}\right)p^{n-k} q^k\]

OpenStudy (e.mccormick):

When k=0, \(p^{n-k}=p^{10}\), right? That is the spot for all to questions randomly being right.

OpenStudy (anonymous):

right.

OpenStudy (e.mccormick):

all to means all ten... don't ask me how it became "to" in my mind. Dyslexia is fun. OK, the questin asks about when 6 are right. So, when is \(p^{n-k}=p^{6}\)?

OpenStudy (anonymous):

when k=4.

OpenStudy (e.mccormick):

Yes! So, we put 4 in there. (By the way: We are answering a and c at the same time.)

OpenStudy (anonymous):

oh sweet! haha so then it should be \[(p+q)^{10}=\sum_{k=4}^{10} (\frac{ 10! }{(6)!4!}) p ^{10}q ^{4}?\]

OpenStudy (e.mccormick):

\[\left[(p+q)^{10}\right]_4=\sum_{k=0}^{10}\left(\frac{10!}{(10-4)!4!}\right)p^{10-4} q^4 \]

OpenStudy (e.mccormick):

Your p also needed to change in the right side... But close.

OpenStudy (e.mccormick):

\[\left[(p+q)^{10}\right]_4=\sum_{k=0}^{10}\left(\frac{10!}{6!4!}\right)p^{6} q^4 \]

OpenStudy (anonymous):

okay so how do I evaluate \[[(p+q)^{10}]_{4}\]

OpenStudy (e.mccormick):

Well, the right ahnd side is what that is equal to. We are saying, with the subscript, take the 4th spot. The right hand side becomes that. Now, have you done the wonderful factorials there?

OpenStudy (anonymous):

working on it

OpenStudy (e.mccormick):

\[\frac{ 10\cdot 9\cdot 8\cdot 7\cdot 6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1}{ (6\cdot 5\cdot 4\cdot 3\cdot 2\cdot 1)(4\cdot 3\cdot 2\cdot 1)} \\ \frac{ 10\cdot 9\cdot 8\cdot 7}{ (4\cdot 3\cdot 2)} \\ 5\cdot 3\cdot 2\cdot 7=210\]

OpenStudy (anonymous):

Yeah I got that too.

OpenStudy (anonymous):

Now what?

OpenStudy (e.mccormick):

got a call... sry... This is taking a bit.

OpenStudy (anonymous):

No worries.(:

OpenStudy (e.mccormick):

Boss at hardware store needed list of parts... OK, so: \(210p^6q^4\) is where we stopped.

OpenStudy (anonymous):

yes.

OpenStudy (e.mccormick):

p is the chance of being correct in a single event, so \(p=\frac{1}{4}\). Now, what was q?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4}\]

OpenStudy (e.mccormick):

So we put those in that formula for that term.

OpenStudy (anonymous):

wait so are we done with the sigme part now?

OpenStudy (anonymous):

sigma

OpenStudy (e.mccormick):

Yah, we used it to find the proper term. That was the goal of sigma. Instead of finding all the terms, we found just the one we needed.

OpenStudy (e.mccormick):

\[210\left(\frac{1}{4}\right)^6\left(\frac{3}{4}\right)^4\]

OpenStudy (anonymous):

alright so p is \[2.44X10^{-4}\]

OpenStudy (anonymous):

and q is 0.316

OpenStudy (anonymous):

right?

OpenStudy (e.mccormick):

The fractions may be easier to deal with. I tend to leave things in fractional form until the last moment. When you become friends with fractions, the chances of you making an error or bad approximation become less.

OpenStudy (anonymous):

okay so p=\[\frac{ 1 }{ 4096 }\]

OpenStudy (anonymous):

and q= \[\frac{ 81 }{ 256 }\]

OpenStudy (e.mccormick):

Yes, which is a power of 2 based number... like 256 is!

OpenStudy (e.mccormick):

\[210\left(\frac{1}{4096}\right)\left(\frac{81}{256}\right)\]

OpenStudy (e.mccormick):

\[\frac{210}{1}\left(\frac{1}{4096}\right)\left(\frac{81}{256}\right)\]\[\frac{105\cdot 2}{1}\left(\frac{1}{4096}\right)\left(\frac{81}{256}\right)\]\[\frac{21\cdot 5 \cdot 2}{1}\left(\frac{1}{4096}\right)\left(\frac{81}{256}\right)\]\[\frac{21\cdot 5 }{1}\left(\frac{1}{2048}\right)\left(\frac{81}{256}\right)\]

OpenStudy (anonymous):

wait, what happened to the 2?

OpenStudy (e.mccormick):

It ran away with the 4096.

OpenStudy (anonymous):

Oh duh!

OpenStudy (e.mccormick):

hehe. Now, because the bottoms are all powers of 2 and the tops no longer have any 2s in them, there is no more similifications.

OpenStudy (anonymous):

so I got .016

OpenStudy (e.mccormick):

\[\frac{8505}{524288}\approx 0.0162220001220703125\]

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

So part a is done right?

OpenStudy (e.mccormick):

And c. Well, c is describing what we did.

OpenStudy (e.mccormick):

Depending on if the teacher wants an exact answer, decimal to so many places, or precentage chance, you will need to pick a format for the answer, but that is all there in the numbers.

OpenStudy (anonymous):

Okay now what about b. I am so thankful for all your help by the way.

OpenStudy (e.mccormick):

Now, the key to "b.Is getting exactly 10 questions correct the same probability as getting exactly zero correct?" is in those fractions and the rules of the Pascal's Triangle. What are the first and last parts of pascals triangle, always, 100% of the time, never any other values.

OpenStudy (e.mccormick):

This is also exposed in choose 10 of 10 \(=\frac{10!}{10!}\)

OpenStudy (anonymous):

The answer is yes right?

OpenStudy (e.mccormick):

We are not there yet. Hehe. The start and end of Pascal's are always 1, right? Just that far at this point.

OpenStudy (anonymous):

okay haha right.

OpenStudy (e.mccormick):

So we have 10 right being \(1p^{10}q^0\) and 0 right being \(1p^{0}q^{10}\) So really, if \(p=q\) they would be the same.

OpenStudy (e.mccormick):

Does \(p=q\)?

OpenStudy (anonymous):

no.

OpenStudy (e.mccormick):

Chance of all 10 right random guessing is:\[\frac{1}{1048576}\] Chance of all 10 wrong random guessing is:\[\frac{59049}{1048576}\] So guessing is really, really bad.

OpenStudy (e.mccormick):

Yah, if chances are equal, the curve is pretty evenly shaped. But when chances get higher for one thing or another, the results skew a lot.

OpenStudy (anonymous):

So the probability of getting 10 correct is not the same as getting zero correct.

OpenStudy (e.mccormick):

Exactly. No where near the same. In fact, remember our 6/10 thing? where we got 0.0162220001220703125 as the probability? Well, 0/10 is 0.05631351470947265625. That is a bigger number!

OpenStudy (e.mccormick):

You have a better chance of getting none correct than of getting 6!

OpenStudy (anonymous):

Hahah holy cow. Thank you so so so so much!(:

OpenStudy (e.mccormick):

Yah, it is almsot 3.5 times more likely that you get 0 right vs. 6. So guessing is really a bad idea.

OpenStudy (e.mccormick):

And it is 59049 times more likely that you get none right vs all 10..... super bad odds there. LOL. Well, have fun! I think you can go back over this and pull out how the Binomial Theorem and factorials lead to a single chance equation, then use that for the explanation in c.